for a high school production of west side story, student tickets cost $5 each whole non-student tickets cost $8. if 480 tickets were sold for the saturday night show and a total of $2895 was collected, how mant tickets of each type were sold?
Answer by stanbon(75887)

For a high school production of west side story, student tickets cost $5 each whole non-student tickets cost $8. if 480 tickets were sold for the saturday night show and a total of $2895 was collected, how many tickets of each type were sold?
------------------------------
Equations:
Quantity Eq.::: x + y = 480
Value Eq.::::: 5x +8y = 2895
------
Multiply thru the Quantity equation by 5
5x + 5y = 5*480
5x + 8y = 2895
---
Subtract the 1st from the 2nd:
3y = 495
y = 165 (# of non-student tickets sold)
---
Since x+y = 480, x = 480-165 = 315 (# of student tickets sold)
--------------------------
Cheers,
Stan H.